Iodine-131 is often used in nuclear medicine to obtain images of the thyroid. If you start with 5.0×10*10 I-131 atoms, how many are left after approximately 1 month? I-131 has a half-life of 8.0 days.
Answer #1

The first order reaction constant is calculated:

k = 0.693 / t1 / 2 = 0.693 / 8 days = 0.087 day-1

The amount of atoms is calculated after 30 days:

ln A = - k * t + ln Ao

ln A = - 0.087 * 30 + ln 5x10 ^ 10

A = 3.7x10 ^ 9 atoms are cleared

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.